OpenStudy (anonymous):

square footage has gone from 1500 sqr. ft to 2330 sqr. ft from 1970 to 2004 Using the model A= Pe^rt r = .008 According to this model which year will the average home exceed 3200 sqr. ft

OpenStudy (campbell_st):

well using the model Initial footage P = 1500 r = 0.008 and A = 3200 you need to find t \[3200 = 1500 e^{0.008t}\] \[e^{0.008t} = 3200/1500\] take the ln of both sides \[0.008t = \ln(3200/1500)\] then \[t = \ln(3200/1500)/0.008\]

OpenStudy (anonymous):

the answer is the year 2028 but i cant figure out how to get to the answer :( ugh so confused

OpenStudy (campbell_st):

once you find t... add it the value to the starting year 1970

OpenStudy (anonymous):

hmm, when i do the equation i get 95? so added to 1970 i get 2065....

OpenStudy (campbell_st):

make sense

OpenStudy (anonymous):

but the real answer is 2028 so no it doesnt make sense .....:(

OpenStudy (campbell_st):

well the problem with your question is your r value... its not 0.008. you need to find r 1st \[2330 = 1500e^{r \times 34}\] solving for r will give r = 0.01295 using this value for r rather than 0.008 and the method above you'll find t = 58